OpenStudy (anonymous):

[2.05] The table below shows the ticket rates for whale watching trips offered by Orca Magic Tours. Age Ticket Price Under three years free Three to twelve years $32 Over twelve years $44 On a certain day, the company took 82 people on whale watching trips. There were 54 children aged twelve and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (compassionate):

What type of equation? Is it linear or a system of equations.

OpenStudy (anonymous):

it mean between which one ogf theese equations (54 + x)32 + (82 – (54 + x)) ∙ 44 = C (54 – x)32 + (82 – 54) ∙ 44 = C (82 + x)32 + (82 – (54 + x)) ∙ 44 = C (82 - x)32 + (82 – 54) ∙ 44 = C

OpenStudy (compassionate):

I'll get back to you in a second. I promise. I am currently helping someone else.

OpenStudy (anonymous):

ok np